OpenStudy (annie96):

1.)Which equation will result if the substitution method is used to solve the following system of linear equations? x+y=6 x=y+2 x+y=y+2 x+(y+2)=6 x=(x-6)+2 (y+2)+y=6 2.)Solve the following system of equations by linear combination: 5x+2y=14 y=x-7 (-3,4) (2,-5) (4,-3) (-5,2)

OpenStudy (anonymous):

Simple substitute x = y + 2 in place of y in first equation and thats ur answer And in second question Just substitute the given values in the answers and see for which values of x and y the given equations holds true and thats ur answer

OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x – y = 7 x = y + 3 @CGGURUMANJUNATH Can you please help me?